OpenStudy (anonymous):

21. d2 + 18d + 81 Answer? what is the proccess???

OpenStudy (ddcamp):

Given: d² + 18*d + 18 = (d+a)(d+b) Solve the system of equations: a + b = 18 a * b = 81

OpenStudy (anonymous):

thank you